# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v1

@@ 4.2. Quantitative Results
All-in-one Image Restoration. We compare our method against five state-of-the-art image restoration approaches, including both CNN-based and transformer-based methods. As shown in Table 1, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets compared to existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, outperforming AdaIR by 0.43dB and 0.011, respectively. We achieve the best performance across all seven sub-tasks. Experimental results indicate that multi-task learning often encounters challenges in balancing performance across different tasks. For instance, AdaIR demonstrates strong performance in deraining but performs poorly in dehazing. On the other hand, OneRestore excels in dehazing but delivers suboptimal results in tasks such as deraining and deblurring. In contrast, our method effectively tackles the balance issue in multi-task learning, achieving consistently optimal performance across all sub-tasks. These findings highlight the effectiveness and superiority of our proposed method.
@@
Degradation-Specific Image Restoration. We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches (Table 2). On Rain100H (Yang et al., 2017), our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing Restormer by 0.84dB and 0.063, respectively. On the LOL dataset (Wei et al., 2018), it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022. For deblurring on GoPro (Nah et al., 2017a), it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003. On RESIDE6K (Qin et al., 2020), it achieves 30.89dB PSNR and 0.973 SSIM, outperforming DehazeFormer by 0.60dB and 0.009. These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
4.3. Qualitative Results
@@
Figure 6 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting. Our method demonstrates superior visual performance compared to other approaches. Specifically, in deblurring, our method restores sharper edge details; in deraining, it more effectively removes rain streaks while recovering background information; and in dehazing, it focuses on regions often overlooked by other methods, resulting in more comprehensive dehazing performance. Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
4.4. Ablation Study
Table 3 presents the results of our ablation study on DPMambaIR for the All-in-One image restoration task, illustrating the impact of removing the degradation prompt, different modulation modes, and the High-frequency Enhancement Block (HEB) from the model. For the baseline model without the Degradation-Aware Prompt Selective Scan and HEB, the PSNR and SSIM are only 26.96 dB and 0.884, respectively. Introducing degradation-awareness via prompts applied to Matrices $A$ and $B$ improves PSNR and SSIM by 0.64 dB and 0.008, respectively, by enabling degradation-aware mechanisms in historical information compression and input feature processing. Adding Prompt $C$, which supplements global degradation information, further increases PSNR by 0.07 dB. The inclusion of HEB significantly boosts performance, preserving high-frequency details. The complete model achieves 27.69 dB PSNR and 0.893 SSIM, demonstrating the importance of all components for optimal performance. Table 4 presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: explicit category labels, a pre-trained degradation extractor from OneRestore (Guo et al., 2024a), and our fine-grained degradation representation via image reconstruction. The OneRestore extractor, designed for classification tasks, performs worse than explicit category labels. In contrast, our method achieves a 0.11 dB PSNR improvement over non-blind restoration with category labels by leveraging fine-grained degradation representations. For degradation utilization, we compare concatenation and attention, with our method outperforming both, demonstrating its effectiveness and superiority in extracting and utilizing degradation information.
For the ablation study on global information supplementation, we compare our method with Mamba’s dual-sequence scanning scheme. As shown in Table 5, our approach not only achieves superior performance compared to the dual-sequence scanning method but also demonstrates a significant reduction in both the Parameters and FLOPs.
Table 3. Ablation Studies on our proposed core Modules.
[…]
@@ 4.4. Ablation Study
Table 4. Ablation Study on Different Degradation Extraction and Utilization Methods.
@@
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
@@
Table 5. Ablation Study on Different Methods for Global Information Supplementation.
@@
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C 27.69 0.893 29.35 153.3
5. Conclusion
@@
In this paper, we proposed DPMambaIR, a novel All-in-One framework for image restoration that balances high restoration quality and computational efficiency across diverse degradation types. DPMambaIR introduces a Degradation-Aware Prompt State Space Model (DP-SSM), which dynamically integrates fine-grained degradation features into the state-space modeling process. This design enables adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, the High-Frequency Enhancement Block (HEB) reduces high-frequency detail loss caused by task competition, ensuring superior edge and texture restoration. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.

### v2

@@ IV-B Quantitative Results
We achieve the best performance across all seven sub-tasks. Experimental results indicate that multi-task learning often encounters challenges in balancing performance across different tasks. For instance, AdaIR demonstrates strong performance in deraining but performs poorly in dehazing. On the other hand, OneRestore excels in dehazing but delivers suboptimal results in tasks such as deraining and deblurring. In contrast, our method effectively tackles the balance issue in multi-task learning, achieving consistently optimal performance across all sub-tasks. These findings highlight the effectiveness and superiority of our proposed method.
@@
Degradation-Specific Image Restoration. We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches, as shown in Table II. On Rain100H [44], our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing AdaIR by 0.84dB and 0.006, respectively. On the LOL dataset [42], it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022. For deblurring on GoPro [43], it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003. On RESIDE6K [45], it achieves 30.89dB PSNR and 0.973 SSIM, outperforming AdaIR by 0.02dB and 0.002. These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
IV-C Qualitative Results
@@
Fig. 7 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting. Our method demonstrates superior visual performance compared to other approaches. Specifically, in deblurring, our method restores sharper edge details; in deraining, it more effectively removes rain streaks while recovering background information; and in dehazing, it focuses on regions often overlooked by other methods, resulting in more comprehensive dehazing performance. Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
IV-D Ablation Study
We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table IV), degradation extraction and utilization methods (Table V), and the global information supplementation strategy (Table VI).
Table IV shows the effectiveness of our core modules. The High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB PSNR by enhancing local details. Degradation-aware prompts ($P_{\Delta}$, $P_{B}$) bring a more significant gain of 0.64 dB, showing the benefit of adaptive modulation. Adding $P_{C}$ for global cues offers further slight improvement. The full model demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance.
TABLE IV: Ablation Studies on Our Proposed Core Modules.
Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.60 0.892 (c) 27.67 0.893 (d) 27.69 0.893
Figure 8: t-SNE visualization of degradation embeddings.
Table V presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [9]; and (3) our fine-grained degradation representation via image reconstruction. The OneRestore extractor, optimized for classification, is insensitive to degradation strength or location, leading to poorer results. Our method surpasses the non-blind ”Classes” baseline by 0.11 dB PSNR.
Fig. 8 visualizes the t-SNE embeddings of the extracted degradation features. Significant gaps exist between different degradation types, indicating clear separability. Notably, although JPEG and noise share the same content during synthesis, the t-SNE visualization reveals that their separability is primarily driven by degradation information rather than image content.
For degradation utilization, we compare concatenation and attention-based fusion with our prompt modulation design. Our prompt-based modulation outperforms simple concatenation or attention-based fusion, showing better adaptability to local degradations.
@@ IV-D Ablation Study
TABLE V: Ablation Study on Different Degradation Extraction and Utilization Methods.
@@
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
Table VI presents the ablation study on global information supplementation. We compare our Prompt-C with Mamba’s bi-directional scanning scheme. Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37% (Table VI). This highlights its effectiveness as a lightweight global information supplementation method.
@@
TABLE VI: Ablation Study on Different Methods for Global Information Supplementation.
@@
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C (Ours) 27.69 0.893 29.35 153.3
V Conclusion
@@
We propose DPMambaIR, a novel All-in-One framework for image restoration capable of handling diverse degradation types. The core of our framework is a Degradation-Aware Prompt State Space Model (DP-SSM) that leverages a fine-grained degradation extractor. This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, a lightweight High-frequency Enhancement Block (HEB) was introduced to complement the main framework by enhancing high-frequency details restoration with negligible computational overhead. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.